# On the slow phase for fixed-energy Activated Random Walks

### v1

@@ 6 One-block estimate
Still assuming $\ell\geq 1$, for $2\leq k\leq k_{0}$, conditioning successively on $\mathcal{G}_{\tau_{\ell-1}+k-1}$ and $\mathcal{G}_{\tau_{\ell-1}+k-2}$ and using (6.2) and (6.3), we get
~$\tilde{\mathbb{P}}(S(\tau_{\ell-1}+k)=0)\geq(1-e^{-150})^{2}>1-k_{0}^{-1}e^{-48}.$
This concludes the case $\ell\geq 1$. Now let us analyze the case $\ell=0$. We remember that we defined $\tau_{-1}=-1$. When $k=1$, $\mathds{1}_{G_{0}}=0$ trivially and (6.7) is valid. For $\ell=0$ and $k=2$, if the first attempted emission was not successful, then $\mathds{1}_{G_{1}}=0$. The only way the first attempt is successful, $S(1)=1$ and $L(1)=0$ occurs if $S(0)=1$. By (6.2), the probability of $S(0)=1$ and $S(1)=1$ is less than $e^{-150}$. Either way, we conclude that (6.7) holds when $k=1$.
For $\ell=0$ and $2<k\leq k_{0}$ we use (6.2) and (6.3) in the same way as before to conclude that ~$\tilde{\mathbb{P}}(S(k)=0)>1-\frac{1}{k_{0}e^{48}}$. Combining (6.5), (6) and (6.7), we conclude (6.1).
The proof of $(ii)$ for $\ell\geq 1$ is almost identical to the proof of $(i)$. In the case $\ell=0$, we cannot use Lemma 6.2 directly. But, since by hypothesis the initial configuration has holes only at $iK$, we can use Remark 6.3 instead of Lemma 6.2 to show that, for $1\leq k\leq k_{0}$, $\tilde{\mathbb{E}}[\mathds{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathds{1}_{\{S(\tau_{\ell-1}+k)=1\}}]<\frac{1}{k_{0}e^{48}}$ and proceed as in $(i)$.
@@ 7 Estimating the hole drift
This section is devoted to prove Lemmas 6.2 and 6.4.
Let $\delta:=(K(1+\lambda))^{-1}$, and, for each $v\in\mathbb{N}$, $v\leq a$, define
$\tilde{Y}_{v}=\begin{cases}+1,&\text{with probability }\frac{\lambda}{1+\lambda};\\ 0,&\text{with probability }\frac{1/2}{1+\lambda}+\delta;\\ -k,&\text{with probability }\frac{1/2}{1+\lambda}\frac{1}{k(k+1)}\text{ for }k=1,2,\dots,v-1;\\ -v&\text{with probability }\frac{1/2}{1+\lambda}-\delta-\sum_{k=1}^{v-1}\frac{1/2}{1+\lambda}\frac{1}{k(k+1)}.\end{cases}$
Lemma 7.1.
@@
Lemma 7.2.
Let $(\tilde{Y}_{a/3}(i))_{i\in\mathbb{N}}$ be a sequence of i.i.d. random variables with common distribution $\tilde{Y}_{a/3}$. Then
$\mathbb{P}\left(\sum_{i=1}^{a/6}\tilde{Y}_{a/3}(i)\geq-a/6\right)\leq e^{-a}\quad\text{ and }\quad\mathbb{P}\left(\sum_{i=1}^{a/2}\tilde{Y}_{a/3}(i)\geq-2a/3\right)\leq e^{-a}.$
@@
Take $k>0$ fixed. For every $b>k\mathbb{E}[\tilde{Y}_{a/3}]$, as $\tilde{Y}_{a/3}\in[-a/3,1]$, it follows from a classical Hoeffding bound that
$\mathbb{P}\left(\sum_{i=0}^{k}\tilde{Y}_{a/3}(i)\geq b\right)\leq\exp\left(-\frac{2(-b+k\mathbb{E}[\tilde{Y}_{a/3}])^{2}}{k(a/3+1)^{2}}\right).$
@@
By Lemma 7.1, $\mathbb{E}[\tilde{Y}_{a/3}]<-40$. We can then take $b=-a/6$ and $k=a/6$ to obtain the first inequality of the lemma, and $b=-2a/3$, $k=a/2$ to obtain the second one. ∎
@@
We call a step of the toppling procedure when the hot particle starts in the hole and either: $(i)$ falls asleep; $(ii)$ is emitted; or $(iii)$ leaves the hole and returns to it. Let $T_{j}$ be the number of steps taken by the procedure between the $(j-1)$-st and the $j$-th attempted emissions.
Lemma 7.3.
For all $v\in\mathbb{Z}_{N}\cap[0,a]$, $\mathbb{P}(T_{j}>a^{3}\medspace|\medspace H(j-1)=v)<a^{-1}$. Moreover, if $v<a/2$, then $\mathbb{P}(T_{j}<a/2\medspace|\medspace H(j-1)=v)<(2a)^{-1}$.
Proof.
@@
First we note that if the hot particle is at the hole, the probability that the next instruction will be a jump instruction is $(\lambda+1)^{-1}$. Therefore, the probability of an emission is at least $((\lambda+1)(3K/2+a))^{-1}$. The first inequality comes then from the fact that, if $a$ is big enough, $(1-((\lambda+1)(3K/2+a))^{-1})^{a^{3}}<a^{-1}$.
@@
If $v<a/2$, then the first $a/2$ steps of an attempted emission cannot end with a failure. If $T_{j}<a/2$, then the $j$th attempted emission ended in an successful emission. The probability of occurrence of an emission is at most $((\lambda+1)K)^{-1}$, then an union bound gives that the probability of an emission occurring in the first $a/2$ steps is at most $(a/2)((\lambda+1)K)^{-1}\leq(2a)^{-1}$ and the second inequality of the lemma follows. ∎
Proof of Lemma 6.2.
@@
We begin by proving (6.2). It is sufficient to prove that, for every $v\in([0,a/2]\cup\{a\})\cap\mathbb{Z}_{N}$, $\mathbb{P}(H(j)>a/2\medspace|\medspace H(j-1)=v)<\frac{4}{a}<e^{-150}$. We split in three cases.
@@
First, we consider $v=a$. In this case, by (P8) every site in the block has a carpet particle, and the hot particle is performing a simple random walk. If the hot particle visits every site in $[0,a]$ before being emitted, the hole is reset to position 00. The probability of this occurring is at least $\frac{K/2-a}{K/2}=1-\frac{2}{a}.$
@@
Then, we look at the case $v\leq a/3$. By Lemma 7.3, it is immediate that ~$\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}\geq a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. If $\{H(j)>a/2\}\cap\{T_{j}<a^{3}\}$ occurs, one of the first $a^{3}$ times the hole is at $a/3$ it moves to $a/2$ before returning to the left of $a/3$. This probability is bounded above by the probability that the sum of $a/6$ independent copies of $\tilde{Y}_{a/3}$ is at least 00. Using Lemma 7.2, ~$\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}>a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. The case follows by an union bound.
@@
Finally, we consider $a/3<v\leq a/2$. A failure cannot occur in the first $a/2-1$ steps. The probability that an emission occurs and $T_{j}<a/2$ is bounded above by $1/4a$ by Lemma 7.3. If an emission does not occur in the first $a/2-1$ steps and the hole is always at the right of $a/3$, the movement of the hole is stochastically dominated by $\tilde{Y}_{a/3}$. Thus the probability of $T_{j}<a/2$ and the hole be always at the right of $a/3$ is bounded above by the probability that the sum of $a/6$ copies of $\tilde{Y}_{a/3}$ is at least $-a/6$, and this probability is at most $e^{-a}$ by Lemma 7.2. Once the hole is at or at the left of $a/3$ we proceed as in the previous case. Hence, we get ~$\tilde{\mathbb{P}}(H(j)>a/2\cap\{T_{j}<a/2\}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$, and ~$\tilde{\mathbb{P}}(H(j)>a/2\cap\{a/2\leq T_{j}<a^{3}\}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. By Lemma 7.3, ~$\tilde{\mathbb{P}}(H(j)>a/2\cap\{T_{j}>a^{3}\}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$, and we conclude by an union bound.
@@
It remains to prove (6.3). Again, it is sufficient to condition on $H(j-1)$. If $H(j-1)\in[0,a/2]\cup\{a\}$, we can use 6.2 to bound the probability of $H(j)\in(a/2,a)$. So let us suppose $H(j-1)\in(a/2,a)$. Since when a failure occurs we have $H(j)=a$, we only need to consider the case of an emission. The probability that an emission occurs in the first $a/2$ steps is bounded above by $\frac{1}{2a}$ by the same argument used in the proof of Lemma 7.3. The probability that a emission does not occur and the hole is never in $[0,a/3]$ in the first $a/2$ steps is at most $e^{-a}$ by Lemma 7.2. Once the hole reaches $[0,a/3]$ we can use Lemma 7.2 again as in the proof of (6.2) to conclude that the probability the hole reaches $(a/2,a)$ is at most $e^{-a}$. An union bound finishes the proof. ∎
Proof of Lemma 6.4.
@@
We notice that if an attempted emission ends in a failure, the next hot particle will be emitted. Thus, for each two attempted emissions, at least one will be successful. The probability that such an emission occurs to the left is at least $\frac{((i+1)K/2)-(iK+a)}{((i+1)K/2)-(iK-3K/2+1)}=\frac{1}{4}-\frac{1}{\sqrt{K}}\geq\frac{1}{5}$. ∎

### v2

@@ 7 One-block estimate
Still assuming $\ell\geq 1$, for $2\leq k\leq k_{0}$, we have that
$\tilde{\mathbb{P}}(S(\tau_{\ell-1}+k)=0)$ $\geq$ $\tilde{\mathbb{E}}[\mathbb{E}[\mathbf{1}_{\{H(\tau_{\ell-1}+k)\in[0,a/2]\}}]\mid\mathcal{G}_{\tau_{\ell-1}+k-1}]$
$\geq$ $\tilde{\mathbb{E}}[(1-e^{-150})\mathbf{1}_{\{H(\tau_{\ell-1}+k-1)\in[0,a/2]\}}]$
$=$ $(1-e^{-150})\tilde{\mathbb{E}}[\mathbb{E}[\mathbf{1}_{\{H(\tau_{\ell-1}+k-1)\in[0,a/2]\}}\mid\mathcal{G}_{\tau_{\ell-1}+k-2}]]$
$\geq$ $(1-e^{-150})^{2}>1-\frac{1}{k_{0}e^{48}},$
where in the second line we used (1) and in the last line we used (4).
This concludes the case $\ell\geq 1$. Now let us analyze the case $\ell=0$. We remember that we defined $\tau_{-1}=-1$. When $k=1$, $\mathbf{1}_{G_{0}}=0$ trivially and (8) is valid. For $\ell=0$ and $k=2$, if the first attempted emission was not successful, then $\mathbf{1}_{G_{1}}=0$. The only way the first attempt is successful, $S(1)=1$ and $L(1)=0$ occurs if $S(0)=1$. By (3), the probability of $S(0)=1$ and $S(1)=1$ is less than $e^{-150}$. Either way, we conclude that (8) holds when $k=1$.
For $\ell=0$ and $2<k\leq k_{0}$ we use (3) and (4) in the same way as before to conclude that $\tilde{\mathbb{P}}(S(k)=0)>1-\frac{1}{k_{0}e^{48}}$. Combining (6), (7) and (8), we conclude (2).
The proof of $(ii)$ for $\ell\geq 1$ is almost identical to the proof of $(i)$. In the case $\ell=0$, we cannot use Lemma 1 directly. But, since by hypothesis the initial configuration has holes only at $iK$, we can use Remark 1 instead of Lemma 1 to show that, for $1\leq k\leq k_{0}$, $\tilde{\mathbb{E}}[\mathbf{1}_{\{L(\tau_{\ell-1}+k)=\ell\}}\mathbf{1}_{\{S(\tau_{\ell-1}+k)=1\}}]<\frac{1}{k_{0}e^{48}}$ and proceed as in $(i)$.
@@ 8 Estimating the hole drift
This section is devoted to prove Lemmas 1 and 2.
The movement of the hole has a drift to the right, at least if it is in the interval $[iK,iK+a/3]$, if $\lambda$ is large. We then choose $a$ large enough such that the hole has a drift to the left as soon as it is at the right of $iK+a/3$. Then, as long as there is no emission, the hole will be “trapped” between those two drifts, and would take an exponentially large number of trials to reach $iK+a$. On the other hand, the event of no emission occurring after an exponential number of trials is unlikely, because the probability of an emission is roughly $1/(\lambda a^{2})$, and this is of order $1/a^{2}$ if $a$ is much larger than $\lambda$.
The probability of occurrence of an emission to the left (or to the right) is bounded above by
$\delta:=\left(\frac{1}{2}\right)\left(\frac{1}{1+\lambda}\right)\left(\frac{1}{K-a}\right).$
For each $v\in\mathbb{N}$, $v\leq a$, define
[…]
Those random variables dominate stochastically the change of the position of the hole, when it is at $iK+v$, conditioned to the event that there are no defects inside the block.
Lemma 3.
@@
Thus, the hole has a drift to the left, when $v\geq a/3$, at least when no emission is made. This is because the excursion of a simple random walk has a heavy tail, and, if we take $a$ large enough, at least one toppling of the hot particle will result in an excursion instead of the particle going to sleep.
Lemma 4.
Let $(\tilde{Y}_{a/3}(i))_{i\in\mathbb{N}}$ be a sequence of i.i.d. random variables with common distribution $\tilde{Y}_{a/3}$. Then there is a constant $\alpha>0$ such that
$\mathbb{P}\left(\sum_{i=1}^{a/6}\tilde{Y}_{a/3}(i)\geq-a/6\right)\leq e^{-\alpha a}\qquad\text{ and }\qquad\mathbb{P}\left(\sum_{i=1}^{a/2}\tilde{Y}_{a/3}(i)\geq-2a/3\right)\leq e^{-\alpha a}.$
@@
Let $(Y^{i})_{i=1}^{+\infty}$ be a sequence of bounded i.i.d. random variables with expectation $\nu<0$. Fix $b>0$. By Chernoff’s concentration inequality, for every $\gamma>0$ such that $\gamma b$ is an integer with $\nu<-\gamma^{-1}$, there is a constant $\alpha^{\prime}>0$ such that
$\mathbb{P}\left(\sum_{i=0}^{\gamma b}Y^{i}>-b\right)\leq e^{-\alpha^{\prime}\gamma b}.$
@@
By Lemma 3, $\mathbb{E}[\tilde{Y}_{a/3}]<-40$. We can then take $b=a/6$ and $\gamma=1$ to obtain the first inequality of the lemma, and b=2a/3b=-2a/3, $\gamma=3/4$ to obtain the second one. ∎
@@
We call a step of the toppling procedure when the hot particle starts in the hole and either: $(i)$ falls asleep; $(ii)$ is emitted; or $(iii)$ leaves the hole and returns to it. Let $T_{j}$ be the number of steps taken by the procedure between the (j1)(j-1)-st and the $j$-th attempted emissions.
Lemma 5.
If $a$ is large enough, for all $v\in\{0,\dots,a\}$, $\mathbb{P}(T_{j}>a^{3}\medspace|\medspace H(j-1)=v)<a^{-1}$. Moreover, if $v<a/2$, then $\mathbb{P}(T_{j}<a/2\medspace|\medspace H(j-1)=v)<a^{-1}$.
The exact value of the constants $1/3$ and $1/2$ is not important. We need the “gap between” $a/3$ and $a/2$ to be able to use the drift to the left of the hole and still have a good number of trials before the hole be able to reach $iK+a$.
Proof of Lemma 5.
@@
First we note that if the hot particle is at the hole, the probability that the next instruction will be a jump instruction is $(\lambda+1)^{-1}$. Therefore, the probability of an emission is at least $((\lambda+1)(3K/2+a))^{-1}$. The first inequality comes then from the fact that, if $a:=\sqrt{K}$ is big enough, $(1-((\lambda+1)(3K/2+a))^{-1})^{a^{3}}<a^{-1}$.
@@
If $v<a/2$, then the first $a/2$ steps of an attempted emission cannot end with a failure. If $T_{j}<a/2$, then the $j$th attempted emission ended in an successful emission. The probability of occurrence of an emission is at most $((\lambda+1)(K-a))^{-1}$, then an union bound gives that the probability of an emission occurring in the first $a/2$ steps is at most $(a/2)((\lambda+1)K)^{-1}\leq a^{-1}$ and the second inequality of the lemma follows. ∎
Proof of Lemma 1.
@@
We begin by proving (3). It is sufficient to prove that, for every $v\in\{0,a/2\}\cup\{a\}$, $\mathbb{P}(H(j)>a/2\medspace|\medspace H(j-1)=v)<\frac{4}{a}<e^{-150}$. We split in three cases.
@@
First, we consider $v=a$. In this case, by (P8) every site in the block has a carpet or a frozen particle, and the hot particle is performing a simple random walk. If the hot particle visits every site in $[0,a]$ before being emitted, the hole is reset to position 0. The probability of this occurring is at least $\frac{K/2-a}{K/2}=1-\frac{2}{a}.$
@@
Then, we look at the case $v\leq a/3$. By Lemma 5, it is immediate that $\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}\geq a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. If $\{H(j)>a/2\}\cap\{T_{j}<a^{3}\}$ occurs, one of the first $a^{3}$ times the hole is at $a/3$ it moves to $a/2$ before returning to the left of $a/3$. This probability is bounded above by the probability that the sum of $a/6$ independent copies of $\tilde{Y}_{a/3}$ is at least 0. Using Lemma 4, $\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}<a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. The case follows by an union bound.
@@
Finally, we consider $a/3<v\leq a/2$. A failure cannot occur in the first a/21a/2-1 steps. The probability that an emission occurs and $T_{j}<a/2$ is bounded above by $1/a$ by Lemma 5. If an emission does not occur in the first a/21a/2-1 steps and the hole is always at the right of $a/3$, the movement of the hole is stochastically dominated by $\tilde{Y}_{a/3}$. Thus the probability of $T_{j}<a/2$ and the hole be always at the right of $a/3$ is bounded above by the probability that the sum of $a/6$ copies of $\tilde{Y}_{a/3}$ is at least a/6-a/6, and this probability is at most $e^{-\alpha a}$ by Lemma 4. Once the hole is at $a/3$ or to the left of $a/3$ we proceed as in the previous case. Hence, we get $\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}<a/2\medspace|\medspace H(j-1)=v)<\frac{1}{a}$, and $\tilde{\mathbb{P}}(H(j)>a/2\text{ and }a/2\leq T_{j}<a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$. By Lemma 5, $\tilde{\mathbb{P}}(H(j)>a/2\text{ and }T_{j}>a^{3}\medspace|\medspace H(j-1)=v)<\frac{1}{a}$, and we conclude by an union bound.
@@
It remains to prove (4). Again, it is sufficient to condition on H(j1)H(j-1). If $H(j-1)\in[0,a/2]\cup\{a\}$, we can use (3) to bound the probability of $H(j)\in(a/2,a)$. So let us suppose $H(j-1)\in(a/2,a)$. Since when a failure occurs we have $H(j)=a$, we only need to consider the case of an emission. The probability that an emission occurs in the first $a/2$ steps is bounded above by $\frac{1}{a}$ by the same argument used in the proof of Lemma 5. The probability that a successful emission does not occur and the hole is never in $[0,a/3]$ in the first $a/2$ steps is at most $e^{-\alpha a}$ by Lemma 4. Once the hole reaches $[0,a/3]$ we can use Lemma 4 again as in the proof of (3) to conclude that the probability the hole reaches $(a/2,a)$ is at most $e^{-\alpha a}$. An union bound finishes the proof. ∎
Proof of Lemma 2.
@@
We notice that if an attempted emission ends in a failure, the next hot particle will be emitted. Thus, for each two attempted emissions, at least one will be successful. The probability that such an emission occurs to the left is at least $\frac{((i+1)K/2)-(iK+a)}{((i+1)K/2)-(iK-3K/2+1)}=\frac{1}{4}-\frac{1}{\sqrt{K}}\geq\frac{1}{5}.$
Acknowledgment
The authors are grateful to a reviewer for thoughtful and constructive remarks which led to a considerable improvement in the presentation.
Funding Information
B. N. B. L.  was partially supported by CNPq grant 315861/2023-1. L. R.  was partially supported by FAPESP grant 2023/13453-5. C. T.  was partially supported by CAPES grant 88887.595894/2020-00, FAPEMIG grant 15435 and CNPq grant 151432/2024-4.